Pagami Fire Still Hanging On

     The weather I find so beautiful is not so wonderful for the folks working on the Pagami Creek Fire.  The temperatures in the 70’s, clear skies and dry weather makes smoldering surfaces flare up.  The crew is forever looking for hot spots and still trying to contain the remaining 29% of the perimeter.  With wind expected to pick up the crew is shuffling a bit to be prepared.  See below for details.


TODAY’S MESSAGE: The Pagami Creek Fire remains at 92,682 acres and is 71% contained. There may be increased smoke in the area as fire activity increases (especially on the east side) with higher temperatures and wind increase over the fire area. Hot spots outside the fire perimeter continue to be located and contained. Crews are being moved in anticipation of stronger southerly winds expected on Thursday. Additional air resources (two CL215 water scooping aircraft and one type 1 helicopter) are being assigned to the fire and will become available through the day as they arrive.


· On the West and Northwest portions of the fire crews will patrol, mop-up existing lines and back haul unneeded equipment.

· On the Northeast and East portions of the fire, line construction will continue as well as patrol and mop-up of previously constructed line.

· On the Southern portions near Isabella, mop-up and rehabilitation of dozer line continues.

· On the Southwestern portions, line construction will continue between Superstition and Bog Lakes.

· All crews will be checking outside the line for hot spots that are expected to show up with the warmer weather and increased wind.

· Air resources will continue to provide support to firefighters on the ground.

WEATHER: Dry and warm weather with increasing winds is forecast over the fire through the week. High temperature today is expected to be 79oF with winds out of the southeast at 6 to 11 mph with gusts to 15 mph.

AREA/ROAD CLOSURES: The Superior National Forest has reduced the closure area around the Pagami Creek Fire. Some areas that have been reopened include Snowbank Lake, North Kawishiwi River, and Bald Eagle Lake. A small portion of the BWCAW and national forest lands remain closed. To see the updated closure, please visit or

The road closures still in effect form a continual route from west to east. The closure includes the following roads and areas: FR 377D from the intersection of FR 377 (Tomahawk Rd) to the BWCAW boundary (Bog Lake Access); FR 377 (Tomahawk Rd) from the intersection of FR 377D east to FR 379; FR 379 south to the intersection of FR 369; the area north of FR 369 between the intersection of FR 379 and FR 356 (FR 369 is OPEN); FR 356 northeast from the intersection of FR 369 to FR 1244 (Coffee Lake Access); FR 1244 south to the intersection of Lake County Road 7 (CR 7); All areas west of CR 7 from the intersection of FR 1244 to FR 354 (CR 7 is OPEN); FR 354 north to the BWCAW boundary (Kawishiwi Lake Access).

FIRE RESTRICTIONS: The Superior National Forest has lifted all campfire restrictions. For more information, please visit

LOCATION: The fire started approximately 13 miles east of Ely.

CAUSE: Lightning

DETECTED: August 18, 2011

RESOURCES: There are currently 530 personnel assigned to the incident. Resources include: 12 hotshot (Type 1) crews, 2 Type 2 crews, 5 engines, 0 dozers, 2 water tenders, 2 camp crews and other personnel. The following air resources are available: 3 Type 1 helicopters, 1 Type 3 helicopters, and three Beaver aircraft on floats.


Area closures and burning restrictions:

Smoke conditions:



Basic Information

Incident Type Wildfire
Cause Lightning
Date of Origin Thursday August 18th, 2011 approx. 12:00 AM
Location Began In the Boundary Waters Canoe Area Wilderness, approximately 14 miles east of Ely.
Incident Commander Greg Peterson

Current Situation

Total Personnel 454
Size 92,682 acres
Percent Contained 71%
Fuels Involved

10 Timber (litter and understory) FM10, TU2

Fire Behavior

Creeping and light surface fire activity with single tree torching.


Planned Actions

Along north, northwest, and southeast fire perimeter, Division A/B/C/W/X/Z: Patrol and Mop up. On the northeast and eastern flank in Division D/E/T/U/V/Y: Continue direct line construction, hold and improve existing line. All Divisions will monitor for spots outside the line. Rehab Group: Continue rehab of dozer lines.

Growth Potential


Terrain Difficulty



Peterson’s incoming IMT assumed command of the fire Monday at 0600.

Current Weather

Wind Conditions 6-11 mph S
Temperature 79 degrees
Humidity 33%